# MANet: Fine-Tuning Segment Anything Model for Multimodal Remote Sensing Semantic Segmentation

### v1

@@ III-A The Standard Adapter
$\bm{x}_{o}$ $=$ $\mathrm{MLP}\left(\mathrm{LN}(\bm{x}_{i})\right)+s\cdot\bm{x}_{a}+\bm{x}_{i},$ (2)
where $s$ is a scaling factor to weight the task-specific and task-agnostic knowledge.
@@
Since the Adapter proposed in [41] was designed for single-modal data, it is referred to as the standard Adapter in the sequel.
III-B Multimodal Adapter (MMAdapter)
@@
Next, we extend the standard Adapter to multimodal tasks. The proposed MMAdapter is the core multimodal fine-tuning technology of this work. As illustrated in Fig. 3(a), we employ dual branches with shared weights to process multimodal information. The Adapter after the Multi-head Attention is retained to extract the features of each modality independently, while the Adapter during the MLP stage is replaced with the proposed MMAdapter. The detailed structure of MMAdapter is shown in Fig. 3(b). While preserving the core structure of the Adapter, the MMAdapter enables modality interaction through the introduction of two weighting coefficients, $\lambda_{1}$ and $\lambda_{2}$. For two specific multimodal input features, $\bm{x}_{i}\in\mathbb{R}^{h\times w\times c}$ and $\bm{y}_{i}\in\mathbb{R}^{h\times w\times c}$, the process of generating the adapted features with MMAdapter can be described as:
$\bm{x}_{a}$ $=$ $\mathrm{ReLU}\left(\mathrm{LN}(\bm{x}_{n})\cdot\bm{W}_{dx}\right)\cdot\bm{W}_{ux},$ (3)
$\bm{y}_{a}$ $=$ $\mathrm{ReLU}\left(\mathrm{LN}(\bm{y}_{n})\cdot\bm{W}_{dy}\right)\cdot\bm{W}_{uy},$ (4)
@@
where $\bm{W}_{dx},\bm{W}_{dy}\in\mathbb{R}^{c\times\hat{c}}$ and $\bm{W}_{ux},\bm{W}_{uy}\in\mathbb{R}^{\hat{c}\times c}$ are the down projections and up projections, respectively. After that, the multimodal output features $\bm{x}_{o}$ and $\bm{y}_{o}$ are generated using $\lambda_{1}$ and $\lambda_{2}$ as follows:
$\bm{x}_{o}$ $=$ $\mathrm{MLP}\left(\mathrm{LN}(\bm{x}_{i})\right)+\lambda_{1}\cdot\bm{x}_{a}+(1-\lambda_{1})\cdot\bm{y}_{a}+\bm{x}_{i},$ (5)
$\bm{y}_{o}$ $=$ $\mathrm{MLP}\left(\mathrm{LN}(\bm{y}_{i})\right)+\lambda_{2}\cdot\bm{y}_{a}+(1-\lambda_{2})\cdot\bm{x}_{a}+\bm{y}_{i}.$ (6)
@@
During the fine-tuning stage, only the newly added parameters are optimized, while the other parameters remain fixed. Detailed annotations are provided in Fig. 2(b), while other subfigures omit these annotations for clarity and conciseness. Finally, it is worth mentioning that the design shown in Fig. 3 can be generalized to more than two modalities in a straightforward manner.
III-C The Proposed MANet
@@
Fig. 4 shows an overview of the proposed MANet. The input to MANet is first processed by SAM’s image encoder endowed with MMAdapter, responsible for extracting and fusing multimodal remote sensing features using the multimodal fine-tuning mechanism. The output is then fed into DFM, which receives two single-scale multimodal outputs from the encoder and expands them into two sets of multiscale multimodal features using pyramid modules. These high-level abstract features are then fused by four squeeze-and-excitation (SE) Fusion modules to generate a group of multiscale features. Finally, the outputs of DFM are passed to the decoder to produce segmentation prediction maps. In this section, we will elaborate on the key components of the proposed MANet.
@@
  • SAM’s image encoder: We denote the optical images and their corresponding DSM data as $\bm{X}\in\mathbb{R}^{H\times W\times 3}$ and $\bm{Y}\in\mathbb{R}^{H\times W\times 1}$, respectively, where $H$ and $W$ represent the height and width of the inputs. The SAM’s image encoder, employing a non-hierarchical ViT backbone, first embeds the input into a size of $\mathbb{R}^{h\times w\times c}$, where $h=\frac{H}{16}$, $w=\frac{W}{16}$, and $c$ is the embedding dimension. Then stacked ViT Blocks is used to extract features. This feature size is maintained throughout the encoding process [42]. As illustrated in Fig. 4, both $\bm{X}$ and $\bm{Y}$ are input into the SAM’s image encoder. It is important to note that the same SAM encoder is used for DSM data, which demonstrates that SAM can be utilized to extract features from non-image data. The SAM’s image encoder extracts and fuses multimodal features, generating high-level abstract features $\bm{F}_{x}\in\mathbb{R}^{h\times w\times c}$ and $\bm{F}_{y}\in\mathbb{R}^{h\times w\times c}$ through the multimodal fine-tuning modules.
    Figure 5: (a) Schematic of the DFM. There are two Pyramid modules that expand multimodal features for multiscale features before they are fused by four SE Fusion modules. (b) Schematic of the SE Fusion module [16]. Notably, we just employ the existing simple fusion module and do not design this structure specifically, which proves that our primary gains are derived from the multimodal fine-tuning strategy based on the vision foundation model.
@@
  • DFM: Multiscale features play a critical role in semantic segmentation tasks since dense predicting requires both global and local information. As shown in Fig. 5(a), two pyramid modules, each consisting of a set of parallel convolutions or deconvolutions, are used to generate multiscale feature maps. Starting with the plain ViT feature map at a scale of $\frac{1}{16}$, we produce feature maps at scales of $\{\frac{1}{4},\frac{1}{8},\frac{1}{16},\frac{1}{32}\}$ using convolutions with strides of $\{\frac{1}{4},\frac{1}{2},1,2\}$, where fractional strides indicate deconvolutions [42]. These simple pyramid modules generate two sets of multimodal multiscale features, denoted as $\bm{F}_{x}^{i}$ and $\bm{F}_{y}^{i}$, where $i=\{1,2,3,4\}$ represents the scale index. Subsequently, four SE Fusion modules [16] are employed to further integrate the multimodal features. It is worth mentioning that more advanced fusion modules can yield further improved segmentation performance.
@@
    As illustrated in Fig. 5(b), the SE Fusion module begins by aggregating the global information from the multimodal features. For the $i$-th fusion module, with an input channel size of $C_{i}$, the squeeze and excitation process is performed through two convolutional operations with a kernel size of $1\times 1$, followed by the ReLU and Sigmoid activations. The multimodal outputs are then weighted and combined element-wise, producing the enhanced fused features denoted by $\bm{F}_{f}^{i}$. The outputs from the four SE Fusion modules form the multiscale fusion features, denoted as $\bm{F}_{f}^{\emph{1-4}}$, which are fed into the decoder for processing. The decoder introduced in UNetformer [29] is employed in this work that reconstructs abstract semantic information into a segmented map by focusing on both global and local information.
Figure 6: Here we present four data samples of size $2048\times 2048$ from Vaihingen (first two columns) and Potsdam (last two columns), respectively. The first row displays the orthophotos with NIRRG channels for Vaihingen and RGB channels for Potsdam. The second and third rows present the corresponding pixel-wise depth information and ground truth labels. They show the individual and complementary characteristics of remote sensing data from different sources.
TABLE II: Quantitative results on the Vaihingen dataset. The best results are in Bold. The second best results are underlined. (%).
Method Backbone OA mF1 mIoU
Bui. Tre. Low. Car Imp. Total
FuseNet [9] VGG16 96.28 90.28 78.98 81.37 91.66 90.51 87.71 78.71
vFuseNet [30] VGG16 95.92 91.36 77.64 76.06 91.85 90.49 87.89 78.92
ABCNet [43] ResNet18 94.10 90.81 78.53 64.12 89.70 89.25 85.34 75.20
MAResU-Net [44] ResNet18 94.84 89.99 79.09 85.89 92.19 90.17 88.54 79.89
ESANet [45] ResNet34 95.69 90.50 77.16 85.46 91.39 90.61 88.18 79.42
CMGFNet [46] ResNet34 97.75 91.60 80.03 87.28 92.35 91.72 90.00 82.26
PSPNet [47] ResNet101 94.52 90.17 78.84 79.22 92.03 89.94 86.55 76.96
SA-GATE [48] ResNet101 94.84 92.56 81.29 87.79 91.69 91.10 89.81 81.27
CMFNet [14] VGG16 97.17 90.82 80.37 85.47 92.36 91.40 89.48 81.44
UNetFormer [29] ResNet18 96.23 91.85 79.95 86.99 91.85 91.17 89.48 81.97
MFTransNet [15] ResNet34 96.41 91.48 80.09 86.52 92.11 91.22 89.62 81.61
TransUNet [49] R50-ViT-B 96.48 92.77 76.14 69.56 91.66 90.96 87.34 78.26
FTransUNet [16] R50-ViT-B 98.20 91.94 81.49 91.27 93.01 92.40 91.21 84.23
RS3Mamba [50] R18-Mamba-T 97.40 92.14 79.56 88.15 92.19 91.64 90.34 82.78
MANet ViT-B 97.90 91.66 80.98 87.08 93.04 92.12 90.29 82.71
ViT-L 98.84 93.17 81.16 89.23 93.39 92.93 91.51 84.72
ViT-H 98.38 93.94 80.70 90.47 93.59 92.97 91.71 85.03
TABLE III: Quantitative results on the Potsdam dataset. The best results are in Bold. The second best results are underlined. (%).
Method Backbone OA mF1 mIoU
Bui. Tre. Low. Car Imp. Total
FuseNet [9] VGG16 97.48 85.14 87.31 96.10 92.64 90.58 91.60 84.86
vFuseNet [30] VGG16 97.23 84.29 89.03 95.49 91.62 90.22 91.26 84.26
ABCNet [43] ResNet18 96.23 78.92 86.40 92.92 88.90 87.52 88.14 79.26
MAResU-Net [44] ResNet18 96.82 83.97 87.70 95.88 92.19 89.82 90.86 83.61
ESANet [45] ResNet34 97.10 85.31 87.81 94.08 92.76 89.74 91.22 84.15
CMGFNet [46] ResNet34 97.41 86.80 86.68 95.68 92.60 90.21 91.40 84.53
PSPNet [47] ResNet101 97.03 83.13 85.67 88.81 90.91 88.67 88.92 80.36
SA-GATE [48] ResNet101 96.54 81.18 85.35 96.63 90.77 87.91 90.26 82.53
CMFNet [14] VGG16 97.63 87.40 88.00 95.68 92.84 91.16 92.10 85.63
UNetFormer [29] ResNet18 97.69 86.47 87.93 95.91 92.27 90.65 91.71 85.05
MFTransNet [15] ResNet34 97.37 85.71 86.92 96.05 92.45 89.96 91.11 84.04
TransUNet [49] R50-ViT-B 96.63 82.65 89.98 93.17 91.93 90.01 90.97 83.74
FTransUNet [16] R50-ViT-B 97.78 88.27 88.48 96.31 93.17 91.34 92.41 86.20
RS3Mamba [50] R18-Mamba-T 97.70 86.11 89.53 96.23 91.36 90.49 91.69 85.01
MANet ViT-B 97.93 87.13 87.72 95.68 92.68 90.89 91.79 85.14
ViT-L 98.31 88.78 87.27 96.29 93.69 91.62 92.51 86.37
ViT-H 98.44 87.37 90.36 96.24 93.17 91.71 92.70 86.69
Figure 7: Visualized comparisons on the Vaihingen test set with the size of $1800\times 1800$. (a) NIRRG images, (b) DSM, (c) Ground Truth, (d) MAResU-Net, (e) vFuseNet, (f) FuseNet, (g) ESANet, (h) SA-GATE, (i) CMGFNet, (j) TransUNet, (k) CMFNet, (l) UNetFormer, (m) MFTransNet, (n) FTransUNet, (o) $\rm{RS^{3}Mamba}$, (p) The proposed MANet. Purple boxes are added to highlight the differences.
@@ IV-A Datasets
Potsdam: This dataset is much larger than the Vaihingen dataset, which consists of $24$ high-resolution Orthophotos, each with the size of $6000\times 6000$ pixels. It includes four multispectral bands: Infrared, Red, Green, and Blue (IRRGB), along with a normalized DSM of $5$ cm. The last three bands in this dataset are utilized to diversify our experiments. The $24$ orthophotos are split into $18$ patches for training and $6$ for testing. Using the same sliding window approach, this dataset contains $10368$ training samples and $3456$ test samples.
@@
The Vaihingen and Potsdam datasets classify five main foreground categories: Building (Bui.), Tree (Tre.), Low Vegetation (Low.), Car, and Impervious Surface (Imp.). Additionally, a background class labeled as Clutter contains indistinguishable debris and water surfaces. Visual examples from both datasets are presented in Fig. 6. Notably, the sliding window moves with a smaller step size, and the overlapping areas are averaged to the reduced boundary effect during the test stage.

### v2

@@ III-A Standard Adapter and the Proposed MMAdapter
    $\bm{x}^{SA}_{o}$ $=\mathcal{F}(\bm{x}_{i})+s\cdot\bm{x}^{SA}_{a}+\bm{x}_{i},$ (2)
@@
    where $\mathcal{F}(\cdot)$ denotes the MLP operation and $s$ is a scaling factor to weight the task-specific and task-agnostic knowledge. Since the Adapter proposed in [wu2023medical] was designed for single-modal data, it is referred to as the standard Adapter in the sequel.
    Figure 4: (a) The detailed structure of standard LoRA [hu2021lora] and (b) the proposed MMLoRA. It is observed that MMLoRA adopts the design principles of MMAdapter, which not only reduces the module’s complexity but also highlights the versatility of this strategy.
@@
  • Proposed MMAdapter: Next, we extend the standard Adapter to multimodal tasks. The proposed MMAdapter is a core component in the proposed multimodal fine-tuning framework. As illustrated in Fig. 3(c), we employ dual branches with shared weights to process multimodal information. The Adapter after the Multi-head Attention is retained to extract features from each modality independently, while the Adapter during the MLP stage is replaced with the proposed MMAdapter. The detail of MMAdapter is presented in Fig. 3(d). While preserving the core structure of the Adapter, the MMAdapter enables modality interaction through a fusion module. Notably, this design accommodates arbitrary feature fusion strategies. To emphasize the effectiveness of the multimodal fine-tuning framework, we adopt the simplest method of element-wise addition based on two weighting factors, $\lambda_{1}$ and $\lambda_{2}$. For two specific multimodal input features, $\bm{x}_{i}\in\mathbb{R}^{h\times w\times c}$ and $\bm{y}_{i}\in\mathbb{R}^{h\times w\times c}$, the process of generating the adapted features with MMAdapter can be described as:
    $\bm{x}^{MMA}_{a}$ $=\mathrm{ReLU}\left(\mathrm{LN}(\bm{x}_{i})\cdot\bm{W}_{dx}\right)\cdot\bm{W}_{ux},$ (3)
    $\bm{y}^{MMA}_{a}$ $=\mathrm{ReLU}\left(\mathrm{LN}(\bm{y}_{i})\cdot\bm{W}_{dy}\right)\cdot\bm{W}_{uy},$ (4)
@@
    where $\bm{W}_{dx},\bm{W}_{dy}\in\mathbb{R}^{c\times\hat{c}}$ and $\bm{W}_{ux},\bm{W}_{uy}\in\mathbb{R}^{\hat{c}\times c}$ are the down projections and up projections, respectively. After that, the multimodal output features $\bm{x}^{MMA}_{o}$ and $\bm{y}^{MMA}_{o}$ are generated using $\lambda_{1}$ and $\lambda_{2}$ as follows:
    $\bm{x}^{MMA}_{o}$ $=\mathcal{F}(\bm{x}_{i})+\lambda_{1}\cdot\bm{x}^{MMA}_{a}+(1-\lambda_{1})\cdot\bm{y}^{MMA}_{a}+\bm{x}_{i},$ (5)
    $\bm{y}^{MMA}_{o}$ $=\mathcal{F}(\bm{y}_{i})+\lambda_{2}\cdot\bm{y}^{MMA}_{a}+(1-\lambda_{2})\cdot\bm{x}^{MMA}_{a}+\bm{y}_{i}.$ (6)
@@
    Only newly added parameters are optimized, while other parameters remain fixed during fine-tuning. Detailed annotations are provided in Fig. 3(b), while other subfigures omit these annotations for clarity and conciseness.
III-B Standard LoRA and the Proposed MMLoRA
  • Standard LoRA: Foundation models are composed of numerous dense layers, typically utilizing full-rank matrix multiplication. To adapt these pretrained models for specific tasks, LoRA [hu2021lora] assumes that during the adaptation process, the updates to the weights have a lower “intrinsic rank” [aghajanyan2020intrinsic]. This mechanism can be applied to any linear layer. For a pretrained weight matrix $\bm{W}_{0}\in\mathbb{R}^{d\times d}$, the update is expressed with a low-rank decomposition:
    $\bm{W}_{0}+\Delta\bm{W}=\bm{W}_{0}+\bm{B}\bm{A},$ (7)
    where $\bm{B}\in\mathbb{R}^{d\times r}$, $\bm{A}\in\mathbb{R}^{r\times d}$, and the rank $r\ll d$.
    During training, $\bm{W}_{0}$ remains fixed and does not receive gradient updates, while the trainable parameters are contained in $\bm{A}$ and $\bm{B}$. Given an input $\bm{x}_{i}$, the forward computation for the adapted module is represented as:
    $\bm{x}^{SL}_{o}=(\bm{W}_{0}+\Delta\bm{W})\bm{x}_{i}=\bm{W}_{0}\bm{x}_{i}+\bm{x}^{SL}_{a}.$ (8)
    The matrix $\bm{A}$ is initialized with a random Gaussian distribution, while $\bm{B}$ is initialized to 0, resulting in $\Delta\bm{W}=0$ at the start of training. The architecture of LoRA is depicted in Fig. 4(a). Throughout this work, the single-modal implementation of LoRA is referred to as the standard LoRA.
  • Proposed MMLoRA: Similar to MMAdapter, we extend standard LoRA to handle multimodal tasks. As shown in Fig. 4(b), a dual-branch structure with shared weights is employed to process multimodal information. This design enables the learning of task-specific knowledge both within individual modalities and across modalities, facilitated by the fusion module. Given inputs $\bm{x}_{i}$ and $\bm{y}_{i}$, the process of generating the adapted features using MMLoRA can be described as:
    $\bm{x}^{MML}_{o}$ $=\bm{W}_{x0}\bm{x}_{i}+\lambda_{1}\cdot\bm{x}^{MML}_{a}+(1-\lambda_{1})\bm{y}^{MML}_{a},$ (9)
    $\bm{y}^{MML}_{o}$ $=\bm{W}_{y0}\bm{y}_{i}+\lambda_{2}\cdot\bm{y}^{MML}_{a}+(1-\lambda_{2})\bm{x}^{MML}_{a},$ (10)
    where
    $\bm{x}^{MML}_{a}$ $=\bm{B}_{x}\bm{A}_{x}\bm{x}_{i},$ (11)
    $\bm{y}^{MML}_{a}$ $=\bm{B}_{y}\bm{A}_{y}\bm{y}_{i}.$ (12)
    Finally, it is worth mentioning that the design shown in Fig. 3 and Fig. 4 can be generalized to more than two modalities in a straightforward manner.
Figure 5: (a) Overview of the proposed MFNet consisting of a SAM’s image encoder with multimodal fine-tuning, a DFM and a general Decoder. The structure of (b) ViT block with MMAdapter and (c) ViT block with MMLoRA. These constitute the two distinct architectures of MFNet.
III-C The Proposed MFNet
@@
Fig. 5 shows an overview of the proposed MFNet and two different multimodal fune-tuning strategies. The input to MFNet is first processed by SAM’s image encoder endowed with MMAdapter or MMLoRA, which is responsible for extracting and fusing multimodal remote sensing features using the multimodal fine-tuning mechanism. The output is then fed into DFM, which receives two single-scale multimodal outputs from the encoder and expands them into two sets of multiscale multimodal features using pyramid modules. These high-level abstract features are then fused by four squeeze-and-excitation (SE) Fusion modules to generate a group of multiscale features. Finally, the outputs of DFM are passed to the decoder to produce segmentation prediction maps. In this section, we introduce the key components of the proposed MFNet in detail.
@@
  • SAM’s image encoder: We denote the optical images and their corresponding DSM data as $\bm{X}\in\mathbb{R}^{H\times W\times 3}$ and $\bm{Y}\in\mathbb{R}^{H\times W\times 1}$, respectively, where $H$ and $W$ represent the height and width of the inputs. The SAM’s image encoder, employing a non-hierarchical ViT backbone, first embeds the input into a size of $\mathbb{R}^{h\times w\times c}$, where $h=\frac{H}{16}$, $w=\frac{W}{16}$, and $c$ is the embedding dimension. Next, stacked ViT Blocks are used to extract features whose size is maintained throughout the encoding process [li2022exploring]. As illustrated in Fig. 5(a), both $\bm{X}$ and $\bm{Y}$ are input into the SAM’s image encoder. It is important to note that the same SAM encoder is used for DSM data, which demonstrates that SAM can be utilized to extract features from non-image data. The SAM’s image encoder extracts and fuses multimodal features, generating high-level abstract features $\bm{F}_{x}\in\mathbb{R}^{h\times w\times c}$ and $\bm{F}_{y}\in\mathbb{R}^{h\times w\times c}$ through the multimodal fine-tuning modules.
  • ViT block with MMLoRA: Fig.5(b) depicts the architecture of MMAdapter within the ViT block. Conversely, MMLoRA serves as a multimodal fine-tuning method applied in parallel with linear layers. For greater clarity, the structure of the ViT block incorporating MMLoRA is presented in Fig.5(c). In the multi-head attention module, LoRA module is applied to the $q$ and $v$ projection layers [zhang2023customized]. At this stage, multimodal interaction is excluded to concentrate on capturing task-specific information within each modality. In the subsequent MLP layer, MMLoRA module is applied to the linear layers in MLP, facilitating the fusion of multimodal information.
@@
  • DFM: Multiscale features play a critical role in semantic segmentation tasks since dense predicting requires both global and local information. As shown in Fig. 6(a), two pyramid modules, each consisting of a set of parallel convolutions or deconvolutions, are used to generate multiscale feature maps. Starting with the plain ViT feature map at a scale of $\frac{1}{16}$, we produce feature maps at scales of $\{\frac{1}{4},\frac{1}{8},\frac{1}{16},\frac{1}{32}\}$ using convolutions with strides of $\{\frac{1}{4},\frac{1}{2},1,2\}$, where fractional strides indicate deconvolutions [li2022exploring]. These simple pyramid modules generate two sets of multimodal multiscale features, denoted as $\bm{F}_{x}^{i}$ and $\bm{F}_{y}^{i}$, where $i=\{1,2,3,4\}$ represents the scale index. Subsequently, four SE Fusion modules [ma2024ftransunet] are employed to further integrate the multimodal features. It is worth mentioning that more advanced fusion modules can yield further improved segmentation performance.
    Figure 6: (a) Schematic of the DFM. There are two Pyramid modules that expand multimodal features for multiscale features before they are fused by four SE Fusion modules. (b) Schematic of the SE Fusion module [ma2024ftransunet]. Notably, we just employ the existing simple fusion module and do not design this structure specifically, which proves that our primary gains are derived from the multimodal fine-tuning strategy based on the vision foundation model.
@@
    As illustrated in Fig. 6(b), the SE Fusion module begins by aggregating the global information from the multimodal features. For the $i$-th fusion module, with an input channel size of $C_{i}$, the squeeze and excitation process is performed through two convolutional operations with a kernel size of $1\times 1$, followed by the ReLU and Sigmoid activations. The multimodal outputs are then weighted and combined element-wise, producing the enhanced fused features denoted by $\bm{F}_{f}^{i}$. The outputs from the four SE Fusion modules form the multiscale fusion features, denoted as $\bm{F}_{f}^{\emph{1-4}}$, which are fed into the decoder for processing. The decoder introduced in UNetformer [wang2022unetformer] is employed in this work that reconstructs abstract semantic information into a segmented map by focusing on both global and local information.
@@ IV-A Datasets
Potsdam: This dataset is much larger than the Vaihingen dataset, which consists of $24$ high-resolution Orthophotos, each with the size of $6000\times 6000$ pixels. It includes four multispectral bands: Infrared, Red, Green, and Blue (IRRGB), along with a normalized DSM of $5$ cm. The last three bands (RGB) in this dataset are utilized to diversify our experiments. The $24$ orthophotos are split into $18$ patches for training and $6$ for testing. Using the same sliding window approach, this dataset contains $10368$ training samples and $3456$ test samples.
@@
The Vaihingen and Potsdam datasets classify five main foreground categories: Building (Bui.), Tree (Tre.), Low Vegetation (Low.), Car, and Impervious Surface (Imp.). Additionally, a background class labeled as Clutter contains indistinguishable debris and water surfaces. Notably, the sliding window to collect training samples moves with a smaller step size, and the overlapping areas are averaged to the reduced boundary effect during the test stage.
Figure 7: Here we present (a/b): two samples of size $2048\times 2048$ from Vaihingen, (c/d): two samples of size $2048\times 2048$ from Potsdam (last two columns), and (e/f): two samples of size $256\times 256$ from MMHunan. The first row displays the orthophotos with NIRRG channels for Vaihingen, RGB channels for Potsdam, and RGB channels for MMHunan. The second and third rows present the corresponding pixel-wise depth information and ground truth labels. They show the individual and complementary characteristics of remote sensing data from different sources.
MMHunan: This dataset [li2022dkdfn] differs significantly in spatial resolution from that of the ISPRS datasets, with a value of $10$ meters. It contains $500$ Sentinel-2 image patches, each of size $256\times 256$ pixels, accompanied by corresponding Digital Elevation Data. We selected the Red, Green, and Blue bands to form the visible imagery. The dataset includes annotations for $7$ land cover types: Cropland, Forest, Grassland, Wetland, Water, Unused Land, and Built-up Area.
Visual examples from all three datasets are presented in Fig. 7. The substantial differences in both data characteristics and land cover categories greatly enhance the diversity of our experiments.
TABLE I: Quantitative results on the Vaihingen dataset. The best results are in Bold. The second best results are underlined. (%).
Method Backbone OA mF1 mIoU
Bui. Tre. Low. Car Imp. Total
FuseNet [FuseNet] VGG16 96.28 90.28 78.98 81.37 91.66 90.51 87.71 78.71
vFuseNet [vfusenet] VGG16 95.92 91.36 77.64 76.06 91.85 90.49 87.89 78.92
MAResU-Net [MAResUNet] ResNet18 94.84 89.99 79.09 85.89 92.19 90.17 88.54 79.89
ESANet [ESANet] ResNet34 95.69 90.50 77.16 85.46 91.39 90.61 88.18 79.42
CMGFNet [CMGFNet] ResNet34 97.75 91.60 80.03 87.28 92.35 91.72 90.00 82.26
PSPNet [PSPNet] ResNet101 94.52 90.17 78.84 79.22 92.03 89.94 86.55 76.96
SA-GATE [SAGATE] ResNet101 94.84 92.56 81.29 87.79 91.69 91.10 89.81 81.27
CMFNet [ma2022crossmodal] VGG16 97.17 90.82 80.37 85.47 92.36 91.40 89.48 81.44
UNetFormer [wang2022unetformer] ResNet18 96.23 91.85 79.95 86.99 91.85 91.17 89.48 81.97
MFTransNet [he2023mftransnet] ResNet34 96.41 91.48 80.09 86.52 92.11 91.22 89.62 81.61
TransUNet [Transunet] R50-ViT-B 96.48 92.77 76.14 69.56 91.66 90.96 87.34 78.26
FTransUNet [ma2024ftransunet] R50-ViT-B 98.20 91.94 81.49 91.27 93.01 92.40 91.21 84.23
RS3Mamba [ma2024rs] R18-Mamba-T 97.40 92.14 79.56 88.15 92.19 91.64 90.34 82.78
FTransDeepLab [FTransDeepLab] ResNet101 98.11 93.45 80.35 89.98 93.23 92.61 91.00 83.87
MultiSenseSeg [MultiSenseSeg] Segformer-B2 97.91 93.04 81.58 89.06 93.56 92.73 91.42 84.53
MFNet (MMLoRA) ViT-B 97.83 94.26 77.82 85.43 91.98 91.93 89.89 82.09
ViT-L 96.85 92.89 81.09 89.95 93.28 92.22 91.09 83.96
ViT-H 97.98 92.35 82.96 90.09 93.25 92.73 91.50 84.66
MFNet (MMAdapter) ViT-B 98.73 91.41 83.09 85.63 92.91 92.62 90.60 83.24
ViT-L 98.84 93.17 81.16 89.23 93.39 92.93 91.51 84.72
ViT-H 98.38 93.94 80.70 90.47 93.59 92.97 91.71 85.03
TABLE II: Quantitative results on the Potsdam dataset. The best results are in Bold. The second best results are underlined. (%).
Method Backbone OA mF1 mIoU
Bui. Tre. Low. Car Imp. Total
FuseNet [FuseNet] VGG16 97.48 85.14 87.31 96.10 92.64 90.58 91.60 84.86
vFuseNet [vfusenet] VGG16 97.23 84.29 89.03 95.49 91.62 90.22 91.26 84.26
MAResU-Net [MAResUNet] ResNet18 96.82 83.97 87.70 95.88 92.19 89.82 90.86 83.61
ESANet [ESANet] ResNet34 97.10 85.31 87.81 94.08 92.76 89.74 91.22 84.15
CMGFNet [CMGFNet] ResNet34 97.41 86.80 86.68 95.68 92.60 90.21 91.40 84.53
PSPNet [PSPNet] ResNet101 97.03 83.13 85.67 88.81 90.91 88.67 88.92 80.36
SA-GATE [SAGATE] ResNet101 96.54 81.18 85.35 96.63 90.77 87.91 90.26 82.53
CMFNet [ma2022crossmodal] VGG16 97.63 87.40 88.00 95.68 92.84 91.16 92.10 85.63
UNetFormer [wang2022unetformer] ResNet18 97.69 86.47 87.93 95.91 92.27 90.65 91.71 85.05
MFTransNet [he2023mftransnet] ResNet34 97.37 85.71 86.92 96.05 92.45 89.96 91.11 84.04
TransUNet [Transunet] R50-ViT-B 96.63 82.65 89.98 93.17 91.93 90.01 90.97 83.74
FTransUNet [ma2024ftransunet] R50-ViT-B 97.78 88.27 88.48 96.31 93.17 91.34 92.41 86.20
RS3Mamba [ma2024rs] R18-Mamba-T 97.70 86.11 89.53 96.23 91.36 90.49 91.69 85.01
FTransDeepLab [FTransDeepLab] ResNet101 97.58 85.87 90.08 96.94 92.81 90.97 92.08 85.62
MultiSenseSeg [MultiSenseSeg] Segformer-B2 98.32 87.65 89.54 96.27 92.46 91.30 92.35 86.10
MFNet (MMLoRA) ViT-B 97.60 86.45 87.87 94.39 92.44 90.57 91.48 84.61
ViT-L 97.59 88.57 88.34 96.35 92.68 90.99 92.13 85.71
ViT-H 98.19 87.30 89.89 96.27 92.80 91.43 92.49 86.34
MFNet (MMAdapter) ViT-B 97.93 87.13 87.72 95.68 92.68 90.89 91.79 85.14
ViT-L 98.31 88.78 87.27 96.29 93.69 91.62 92.51 86.37
ViT-H 98.44 87.37 90.36 96.24 93.17 91.71 92.70 86.69
Figure 8: Visualized comparisons on the Vaihingen test set with the size of $512\times 512$. (a) IRRG images, (b) DSM, (c) Ground Truth, (d) CMFNet, (e) FTransUNet, (f) MFTransNet, (g) CMGFNet, (h) FTransDeepLab, (i) MultiSenseSeg, (j) The proposed MFNet. To highlight the differences, some purple boxes are added to all subfigures.
Figure 9: Visualized comparisons on the Potsdam test set with the size of $1024\times 1024$. (a) IRRG images, (b) DSM, (c) Ground Truth, (d) CMFNet, (e) FTransUNet, (f) MFTransNet, (g) CMGFNet, (h) FTransDeepLab, (i) MultiSenseSeg, (j) The proposed MFNet. To highlight the differences, some purple boxes are added to all subfigures.